This year, the number of our male research students has increased by 20% and the number of female research students has decreased by 20%. We now have one student more than before. Which of the following could be the number of research students in my class now?


1
Expert's answer
2021-10-26T03:15:56-0400

Last year's number of boys "B". Last year's number of girls "G".

This year "1.2B" and "0.8G"


"1.2B+0.8G= B+G+1"

"0.2B=1+0.2G"

"B=5+G"

For example "B=10, G=5." The number of research students in my class now is


"10+5+1=16"

For example "B=20, G=15." The number of research students in my class now is


"20+15+1=36"